if a+b=10vand a^2+b^2=58 find the value of a^2+b^2+c^2

Grade:9

1 Answers

Mukesh
6 years ago
Here c is not provided. So we will assume that the first equation is actually a+b+c=0.
Therefore, we will get c=-10.
This gives us, a^2+b^2+c^c = (a^2+b^2) + c^2 = 58 + (-10)^2 = 58 + 100 = 158.
 
So, the answer is 158.